# MARS: Audio Generation via Multi-Channel Autoregression on Spectrograms

### v1

@@ 3.2 Evaluation Metrics
  • Reconstruction errors. We report MSE and MAE, computed on mel-spectrograms extracted from audio samples, as recommended in [11].
  • Sample diversity metrics. The Number of Statistically Different Bins (NDB/$k$) [15] measures sample diversity by clustering the training space and comparing the distribution of test samples to training data. Scores near 0 indicate good coverage of the training distribution, while values close to 1 reveal mode collapse. To further assess diversity and semantic consistency, we also adopt the Pitch Inception Score (PIS) and Instrument Inception Score (IIS) [16], which evaluate whether the reconstructed samples capture the true distribution of pitch classes and instrument classes in NSynth. Following [11], we also report these metrics, although they focus on “symbolic” properties of notes rather than audio quality.
  • Embedding-based metrics. Kernel Inception Distance (KID) [17] measures the similarity between the distributions of reference and reconstructed embeddings. Specifically, we compute PKID and IKID [16] using the same Inception networks employed for PIS and IIS, where the maximum mean discrepancy (MMD) is applied to compare the embedding distributions. We also adopt Fréchet Audio Distance (FAD) [18], a widely adopted metric for perceptual quality, which leverages VGGish embeddings of reference and reconstructed audio. Each embedding set is modeled as a multivariate Gaussian distribution, $\mathcal{N}_{o}$ for the original and $\mathcal{N}_{r}$ for the reconstructed samples, and FAD is defined as the Fréchet distance between these two distributions.
@@
3.3 Experimental Results
Table 1 compares our MARS with NSynth [10], DDSP [12], and DiffWave [4]. MARS achieves the best scores in NDB/$k$, PKID, and IKID, indicating superior sample diversity and fidelity in both pitch and timbre. It also ranks among the top methods for MSE, MAE, and FAD, reflecting accurate reconstruction and high perceptual quality. PIS and IIS, which measure similarity on sound properties rather than audio quality, remain competitive.
The last row reports metrics on generated audio, which are not directly comparable to the reconstruction metrics; nonetheless, the generated samples maintain low reconstruction error and perceptual similarity, confirming that MARS effectively preserves signal characteristics during synthesis.
3.4 Effectiveness of CMX in Maintaining Audio Quality
@@
4 Conclusion
In this work, we introduced MARS (Multi-channel AutoRegression on Spectrograms), a novel framework for audio generation that adapts next-scale autoregression from image synthesis to the spectrogram domain. By introducing the channel multiplexing (CMX) technique, MARS reduces spatial resolution while preserving frequency information, enabling scalable training with a manageable memory footprint. A shared tokenizer across scales further ensures consistent discrete representations, allowing the autoregressive model to refine spectrograms hierarchically.

### v2

@@ IV-B Evaluation Metrics
  • Reconstruction errors. We report Mean-Squared-Error (MSE) and the Mean-Absolute-Error (MAE), computed on mel-spectrograms extracted from audio samples, as recommended in [25].
  • Sample diversity metrics. The number of statistically different bins (NDB/$k$) [17] quantifies sample diversity. It is computed via a Voronoi decomposition of $k$-means clusters in the training space. Test samples are assigned to clusters using $L_{2}$ distance, and a two-sample $t$-test is used to identify statistically different bins. The final score is the fraction of bins that differ significantly. An NDB/$k$ score close to 0 indicates good diversity and coverage of the training distribution, whereas values near 1 suggest mode collapse. To further assess sample diversity and semantic consistency, we adopt the Pitch Inception Score (PIS) and Instrument Inception Score (IIS) [15], which measure how well reconstructed samples match the true distributions of pitch and instrument classes in the NSynth dataset. These metrics are computed analogously to the standard Inception Score [18], but rely on two task-specific backbone networks: one trained for pitch classification and the other for instrument classification. Following [25], we also report these metrics, although they focus on “symbolic” properties of notes rather than audio quality.
  • Embedding-based metrics. Kernel Inception Distance (KID) [2] measures the similarity between the distributions of reference and reconstructed embeddings. Specifically, we compute PKID and IKID [15] using the same Inception networks employed for PIS and IIS, where the maximum mean discrepancy (MMD) is applied to compare the embedding distributions.
    We also adopt Fréchet Audio Distance (FAD) [10], a widely adopted metric for perceptual quality, which leverages VGGish embeddings of reference and reconstructed audio. Each embedding set is modeled as a multivariate Gaussian distribution, $\mathcal{N}_{o}$ for the original and $\mathcal{N}_{r}$ for the reconstructed samples, and FAD is defined as the Fréchet distance between these two distributions.
@@
IV-C Experimental Results
Table I compares our MARS with NSynth [6], DDSP [5], and DiffWave [11]. MARS achieves the best scores in NDB/$k$, PKID, and IKID, indicating superior sample diversity and fidelity in both pitch and timbre. It also ranks among the top methods for MSE, MAE, and FAD, reflecting accurate reconstruction and high perceptual quality. PIS and IIS, which measure similarity on sound properties rather than audio quality, remain competitive.
The last column reports metrics on generated audio, which are not directly comparable to the reconstruction metrics; nonetheless, the generated samples maintain low reconstruction error and perceptual similarity, confirming that MARS effectively preserves signal characteristics during synthesis.
IV-D Effectiveness of CMX in Maintaining Audio Quality
@@
V Conclusion
In this work, we introduced MARS (Multi-channel AutoRegression on Spectrograms), a novel framework for sound sample generation that adapts next-scale autoregression from image synthesis to the spectrogram domain. By introducing the channel multiplexing (CMX) technique, MARS reduces spatial resolution while preserving frequency information, enabling scalable training with a manageable memory footprint. The same principle can be extended beyond audio to images, video, medical scans, or structured time series since CMX provides a compact, lossless representation that scales efficiently, reducing memory and compute requirements while maintaining full data fidelity. Within the MARS framework, using a shared tokenizer across scales further ensures consistent discrete representations, allowing the autoregressive model to refine spectrograms hierarchically.
@@ V Conclusion
Unlike approaches targeting end-to-end generation of complete songs (often at limited perceptual quality) from high-level conditioning, our objective is to generate high-fidelity audio samples that can be directly used as building blocks in professional music production workflows. As future work, we plan to extend our evaluation to more complex and diverse music datasets with richer instrumentation, longer temporal contexts, and higher sampling rates (e.g., 48 kHz). We also intend to investigate applications in the speech domain, where modeling fine temporal detail and long-range dependencies is particularly challenging, to further assess the generality and scalability of MARS and CMX in real-world settings.